OpenStudy (anonymous):

At the start of a basketball game, the referee tosses a ball into the air for the jump ball. The equation that models the height of the ball is h(t) = -16t^2 + 30t + 8. During what time interval is the height of the ball above 10 feet? (Round your answer to the nearest tenth of a second.) a) Between 1.1 and 1.6 seconds b) Between 0.3 and 2.1 seconds c) Between 0.1 and 1.8 seconds

OpenStudy (anonymous):

i guess we have to solve \[-16t^2 + 30t + 8\geq 10\]

OpenStudy (anonymous):

or maybe \[16t^2-30t-8\leq -10\] \[16t^2-30t+2\leq 0\] \[8t^2-15t+1\leq 0\] and now the quadratic formula to find the zeros you answer will be between them

OpenStudy (anonymous):

ooops i made a mistake \[16t^2-30t-2\leq 0\] \[8t^2-15t-1\leq 0\] that is better

OpenStudy (anonymous):

i messed this up somehow, answer is \(.069<t< 1.8\)